The Texas Rustler has total assets of $645,563 and an equity multiplier of 1.22. What is the debt-equity ratio?
Answer #1

Equity multiplier=Total assets/equity

equity=645,563/1.22

=$529150

Total assets=Total liabilities+Total equity

Hence debt=645,563-529150

=116413

Debt to equity ratio=debt/equity

=116413/529150

=0.22